# Differential Privacy as a Perk: Federated Learning over Multiple-Access Fading Channels with a Multi-Antenna Base Station

### v1

@@ III-B User-level Privacy Guarantee
$\mathcal{D}_{\alpha}^{\left((1+\kappa^{(t)})z\right)}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\nu})})$ (20)
@@
$\Phi$ $=\left(\sqrt{\phi_{T-1}}+\frac{(1+\kappa_{\max})\sqrt{r}Dn}{2\eta c\|\boldsymbol{w}^{(T-1)}\|}\right)^{2},$
@@ Appendix A Important Lemmas
where the first step is by the triangle inequality, Assumption 1, and Lemma A.2, and the second step is by the definition of $\boldsymbol{\mu}^{\prime}$. Thus, we have
$\mathcal{D}_{\alpha}^{((1+\kappa^{(t)})z)}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\nu})})$ (63)
@@
$\mathcal{D}_{\alpha}^{((1+\kappa^{(t)})z)}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\nu})})$ (65)
@@
Then, we consider a real sequence $\{a_{k}\}_{k=\tau}^{T-1}$ and any $\tau\in\{0,1,\cdots,T-1\}$ such that $z_{t}=(1+\kappa_{\max})^{t-\tau}\Xi_{\tau}-\sum_{k=\tau}^{t-1}(1+\kappa_{\max})^{t-k-1}a_{k}$ is non-negative for all $t\geq\tau$ and $z_{T}=0$. By this way, we have $z_{\tau}=\Xi_{\tau}$ and $z_{t+1}=(1+\kappa_{\max})z_{t}-a_{t}$. Furthermore, the proof is conducted by induction, utilizing Lemma III.2 and Lemma III.3. Specifically, we have
@@
$=\mathcal{D}_{\alpha}^{((1+\kappa_{\max})z_{t})}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\theta}^{(t)})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\theta}^{\prime(t)})})+\frac{\alpha a_{t}^{2}r^{2}n^{2}}{2(1-\beta^{(t)})W^{(t)}}$
$\leq\mathcal{D}_{\alpha}^{((1+\kappa^{(t)})z_{t})}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\theta}^{(t)})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\theta}^{\prime(t)})})+\frac{\alpha a_{t}^{2}r^{2}n^{2}}{2(1-\beta^{(t)})W^{(t)}}$
@@
Note that $\mathcal{D}_{\alpha}^{(z_{\tau})}(\mathbb{P}_{\boldsymbol{\theta}^{(\tau)}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime(\tau)}})=0$. Let $a_{\tau}=(1+\kappa_{\max})\Xi_{\tau}$ and $a_{t}=0$ for all $t>\tau$, one may obtain
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}^{(T)}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime(T)}})\!\leq\!\sum_{t=\tau}^{T-1}\frac{2\alpha rc^{2}\phi_{t}}{\beta^{(t)}\sigma^{2}}\!+\!\frac{\alpha(1+\kappa_{\max})^{2}\Xi_{\tau}^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(\tau)})\|\boldsymbol{w}^{(\tau)}\|^{2}}$ (70)
@@
$+\frac{\alpha(1+\kappa_{\max})^{2}D^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(\tau)})\|\boldsymbol{w}^{(\tau)}\|^{2}}\biggr\}.$
@@
$\leq\min_{\beta^{(T-1)}}\frac{2\alpha rc^{2}\phi_{T-1}}{\beta^{(T-1)}\sigma^{2}}+\frac{\alpha(1+\kappa_{\max})^{2}D^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(T-1)})\|\boldsymbol{w}^{(T-1)}\|^{2}}$
$\leq\left(\sqrt{\frac{2\alpha rc^{2}\phi_{T-1}}{\sigma^{2}}}+\sqrt{\frac{\alpha(1+\kappa_{\max})^{2}D^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}\|\boldsymbol{w}^{(T-1)}\|^{2}}}\right)^{2}$
@@
$\Phi=\left(\sqrt{\phi_{T-1}}+\frac{(1+\kappa_{\max})\sqrt{r}Dn}{2\eta c\|\boldsymbol{w}^{(T-1)}\|}\right)^{2}.$ (75)

### v2

@@ III-B User-level Privacy Guarantee
$\mathcal{D}_{\alpha}^{\left((1+\kappa^{(t)})^{Q}z\right)}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\nu})})$ (20)
@@
$\Phi$ $=\left(\sqrt{\phi_{T-1}}+\frac{(1+\kappa_{\max})^{Q}\sqrt{r}Dn}{2\eta c\|\boldsymbol{w}^{(T-1)}\|}\right)^{2},$
@@ Appendix A Important Lemmas
where the first step is by the triangle inequality, Assumption 1, and Lemma A.2, and the second step is by the definition of $\boldsymbol{\mu}^{\prime}$. When round $t$ involves $Q$ successive local SGD updates, the overall mapping $\psi^{(t)}$ can be viewed as the composition of $Q$ such one-step mappings. Then, the $Q$-step bound immediately follows as
$\|\psi^{(t)}(\boldsymbol{\mu})-\psi^{(t)}(\boldsymbol{\mu}^{\prime})\|\leq(1+\kappa^{(t)})^{Q}z.$ (63)
Thus, we have
$\mathcal{D}_{\alpha}^{((1+\kappa^{(t)})^{Q}z)}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\nu})})$ (64)
@@
$\mathcal{D}_{\alpha}^{((1+\kappa^{(t)})^{Q}z)}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\mu})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\nu})})$ (66)
@@
Then, we consider a real sequence $\{a_{k}\}_{k=\tau}^{T-1}$ and any $\tau\in\{0,1,\cdots,T-1\}$ such that $z_{t}=((1+\kappa_{\max})^{Q})^{t-\tau}\Xi_{\tau}-\sum_{k=\tau}^{t-1}((1+\kappa_{\max})^{Q})^{t-k-1}a_{k}$ is non-negative for all $t\geq\tau$ and $z_{T}=0$. By this way, we have $z_{\tau}=\Xi_{\tau}$ and $z_{t+1}=(1+\kappa_{\max})^{Q}z_{t}-a_{t}$. Furthermore, the proof is conducted by induction, utilizing Lemma III.2 and Lemma III.3. Specifically, we have
@@
$=\mathcal{D}_{\alpha}^{((1+\kappa_{\max})^{Q}z_{t})}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\theta}^{(t)})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\theta}^{\prime(t)})})+\frac{\alpha a_{t}^{2}r^{2}n^{2}}{2(1-\beta^{(t)})W^{(t)}}$
$\leq\mathcal{D}_{\alpha}^{((1+\kappa^{(t)})^{Q}z_{t})}(\mathbb{P}_{\psi^{(t)}(\boldsymbol{\theta}^{(t)})}||\mathbb{P}_{\psi^{(t)\prime}(\boldsymbol{\theta}^{\prime(t)})})+\frac{\alpha a_{t}^{2}r^{2}n^{2}}{2(1-\beta^{(t)})W^{(t)}}$
@@
Note that $\mathcal{D}_{\alpha}^{(z_{\tau})}(\mathbb{P}_{\boldsymbol{\theta}^{(\tau)}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime(\tau)}})=0$. Let $a_{\tau}=(1+\kappa_{\max})^{Q}\Xi_{\tau}$ and $a_{t}=0$ for all $t>\tau$, one may obtain
$\mathcal{D}_{\alpha}(\mathbb{P}_{\boldsymbol{\theta}^{(T)}}||\mathbb{P}_{\boldsymbol{\theta}^{\prime(T)}})\!\leq\!\sum_{t=\tau}^{T-1}\frac{2\alpha rc^{2}\phi_{t}}{\beta^{(t)}\sigma^{2}}\!+\!\frac{\alpha(1+\kappa_{\max})^{2Q}\Xi_{\tau}^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(\tau)})\|\boldsymbol{w}^{(\tau)}\|^{2}}$ (71)
@@
$+\frac{\alpha(1+\kappa_{\max})^{2Q}D^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(\tau)})\|\boldsymbol{w}^{(\tau)}\|^{2}}\biggr\}.$
@@
$\leq\min_{\beta^{(T-1)}}\frac{2\alpha rc^{2}\phi_{T-1}}{\beta^{(T-1)}\sigma^{2}}+\frac{\alpha(1+\kappa_{\max})^{2Q}D^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}(1-\beta^{(T-1)})\|\boldsymbol{w}^{(T-1)}\|^{2}}$
$\leq\left(\sqrt{\frac{2\alpha rc^{2}\phi_{T-1}}{\sigma^{2}}}+\sqrt{\frac{\alpha(1+\kappa_{\max})^{2Q}D^{2}r^{2}n^{2}}{2\eta^{2}\sigma^{2}\|\boldsymbol{w}^{(T-1)}\|^{2}}}\right)^{2}$
@@
$\Phi=\left(\sqrt{\phi_{T-1}}+\frac{(1+\kappa_{\max})^{Q}\sqrt{r}Dn}{2\eta c\|\boldsymbol{w}^{(T-1)}\|}\right)^{2}.$ (76)